OpenStudy (rosedewittbukater):

Find the derivative of y = cosx / 1+sinx I used the quotient rule using the derivatives of cos and sin but it just turned into one big equation. Can anyone help?

OpenStudy (rosedewittbukater):

@cwrw238 Do you think you could help me with this?

OpenStudy (tkhunny):

Why would you use the quotient rule? \(y = \cos(x) / 1 + \sin(x) = \cos(x) + \sin(x)\) \(y' = -\sin(x) + \cos(x)\) On the other hand, if you remember your Order of Operations, and you REALLY mean: \(y = \cos(x) / (1 + \sin(x)) = \dfrac{\cos(x)}{1+\sin(x)}\), then that would be a Quotient Rule. Show your work. Someone will help you with the calculus, now that the algebra is straightened out.

OpenStudy (tkhunny):

You may also wish to remember your trig identities.

OpenStudy (rosedewittbukater):

How did you go straight to -sinx+cosx ?

OpenStudy (rosedewittbukater):

I was looking at the trig identities but couldn't find anything to fit in the original equation.

OpenStudy (tkhunny):

You are not paying attention. I worked the problem you WROTE. I did not work the problem you INTENDED.

OpenStudy (cwrw238):

[ -sin x ( 1 + sinx) - cos x ( cos x)] / ( 1 + sin x)^2

OpenStudy (rosedewittbukater):

@cwrw238 That's what I did and then expanded the equation and tried to simplify by using the half-angle formulas. Should I just leave it as it is?

OpenStudy (tkhunny):

\(\sin^{2}(x) + \cos^{2}(x) = 1\) Half angle? Why?

OpenStudy (cwrw238):

= [- sin x - sin^2x - cos^2 x ]/ ( 1 + sin x)^2

OpenStudy (cwrw238):

yes you can use sin^2x + cos^2 x = 1 to simplify it further

OpenStudy (rosedewittbukater):

This is what I got after using the quotient rule and half-angle formula. \[y=\frac{ -\sin(x) + \cos(2x) }{ (\frac{ 1-\cos(x) }{ 2 })+2\sin(x) + 1 }\]

OpenStudy (cwrw238):

theres no need to use the half angle formula

OpenStudy (rosedewittbukater):

What do I do then?

OpenStudy (cwrw238):

did you notice - my posts got swopped around weird!!

OpenStudy (rosedewittbukater):

Sorry my internet isn't working right

OpenStudy (cwrw238):

neither is mine..

OpenStudy (rosedewittbukater):

So after the quotient rule, what next? I don't understand how sin^2x + cos^2x can fit in because its negative in the equation.

OpenStudy (cwrw238):

ok (-sin x - sin^2 x - cos^2 x) / (1 + sin x)^2 = - sin x - ( sin^2x + cos^2 x) / ( 1 + sin x)^2

OpenStudy (cwrw238):

now - what i s next?

OpenStudy (rosedewittbukater):

Oh! I see. Next, do I expand the denominator? (1+sinx)^2 ?

OpenStudy (cwrw238):

no just simplify the numerator then you can cancel

OpenStudy (rosedewittbukater):

So (-sinx -1)/(1+sinx)^2

OpenStudy (rosedewittbukater):

Does that become -(1+sinx)/(1+sinx)^2 ?

OpenStudy (cwrw238):

right - now you can further simplify

OpenStudy (cwrw238):

yes

OpenStudy (rosedewittbukater):

And then (-1)/(1+sinx)

OpenStudy (cwrw238):

right

OpenStudy (rosedewittbukater):

I don't think I can go any further, can I?

OpenStudy (cwrw238):

no

OpenStudy (rosedewittbukater):

Ok

OpenStudy (rosedewittbukater):

Thank you so much for your help! I get it now.

OpenStudy (cwrw238):

yw